OpenStudy (anonymous):

A curve is defined by the equations: x = 2+2 cos(t) y = 1-3 sin^2(t) Find an equation in x and y for this curve. I know the answer is y=3/4x^2-3x+1 But how do I get there?

Directrix (directrix):

x = 2+2 cos(t) --> Solve for cos(t) --> 2cos(t) = x - 2 --> cos(t) = (x-2)/2 y = 1-3 sin^2(t) --> Solve for sin^2(t) --> -3sin^2(t) = y - 1 --> sin^2(t) = (y-1)/(-3)

Directrix (directrix):

Recall the trig identity: sin^2(t) + cos^2(t) = 1 Into that, substitute the results from above: cos(t) = (x-2)/2 and sin^2(t) = (y-1)/(-3) Simplify and see what you get. @Loulita

OpenStudy (anonymous):

@Directrix Since cos(t) is not squared, do I have to substitute it as ((x-2)/2)^2?

OpenStudy (anonymous):

yes

Directrix (directrix):

I agree.

OpenStudy (anonymous):

I am having trouble with the simplification. Here is where I am: (y-1)/3+((x-2)/2)^2=1 (y-1)/3+(x^2-2x+4)/4=1 @Directrix

Directrix (directrix):

@Loulita Is this what you are crunching? ((x-2)/2)^2 + ((y-1)/(-3)) = 1

OpenStudy (anonymous):

Yes it is. I may have simplified ((x-2)/2)^2 incorrectly but I'm not sure.

Directrix (directrix):

((x-2)/2)^2 is not equal to (x^2-2x+4)/4. Do you see the error?

Directrix (directrix):

What is (x - 2)^2 ?

Directrix (directrix):

((x-2)/2)^2 + ((y-1)/(-3)) = 1 cranks out to y = (3*x^2)/4 - 3x + 1

OpenStudy (anonymous):

Oh! I knew I was making a silly mistake somewhere. Thank you!